Solve for r. H= 1/2r + 3b -1/4
I have r= 12b-4h-1/2 (everything is over 2 in fraction form) is this right?

Found 2 solutions by Alan3354, mananth:
Answer by Alan3354(69443)

If you mean (1/2)r, then:
H= 1/2r + 3b -1/4
2H = r + 6b - 1/2
r = 2H - 6b + 1/2

Answer by mananth(16949)

H= 1/2r + 3b -1/4 ( iam not sure whether it is (1/2) r OR 1/2r
Assume (1/2 * r)
..
H-3b+1/4 = 1/2 *r
r=2(H-3b+1/4)
r=2H-6b+1/2
..
Assume (1/2r)
H= 1/2r +3b -1/4
H-3b+1/4= 1/2r
2r= 1/(h-3b+1/4)
2r=1/(4h-12b+1)/4)
2r=1*4/(4h-12b+1)
r= 2/(4H-12b+1)